# Fast Robot Adaptation via Hand Path Retrieval

### v3

@@ 3.3 Putting it All Together: Fast-Adaptation with Parameter-Efficient Policy Fine-tuning
Adapting to $\mathcal{D}_{\text{retrieved}}$. To rapidly adapt to a new task with minimal data, we leverage parameter-efficient fine-tuning using task-specific adapters—small trainable modules that modulate the behavior of the frozen base policy. Adapter-based methods have shown promise in few-shot imitation learning [37, 38], making them ideal for our limited retrieved dataset $\mathcal{D}_{\text{retrieved}}$. Following the findings of Liu et al. [38], we specifically insert LoRA layers [39] into the transformer blocks of $\pi_{\text{base}}$. These are low-rank trainable matrices (typically $0.1\%$$2\%$ of the base policy’s parameters) inserted between the attention and feedforward layers (see Figure 2, LoRA Layers). During fine-tuning, we keep $\pi_{\text{base}}$ frozen and update only the parameters of these LoRA layers, $\theta$, using $\mathcal{D}_{\text{retrieved}}$.
Loss Re-Weighting. While our retrieval mechanism identifies sub-trajectories relevant to the target task, not all will be equally useful. To prioritize the most behaviorally aligned examples, we reweight the BC loss with an exponential term $\in(0,\infty)$ following Advantage-Weighted Regression [40], where each sub-trajectory is weighted based on its similarity (from S-DTW) to the hand demonstration. Intuitively, this upweights the loss of the most relevant examples in $\mathcal{D}_{\text{retrieved}}$ and conversely downweights those that are less relevant. Finally, because trajectory cost scales vary depending on the task being retrieved and the features being used for S-DTW, we rescale the S-DTW costs $C_{i,\text{path}}$ to a fixed range. For each $\tau_{i}\in\mathcal{D}_{\text{retrieved}}$, its weight $e^{-C_{i,\text{path}}}$ is scaled to between $[0.01,100]$, where the normalization term comes from the sum of costs of all trajectories in $\mathcal{D}_{\text{retrieved}}$. Our final training loss is:
$\mathcal{L}_{\text{BC};\theta}=\frac{1}{|\mathcal{D}_{\text{retrieved}}|}\sum_{\tau_{i}\in\mathcal{D}_{\text{retrieved}}}\underbrace{\exp(-C_{i,\text{path}})}_{\text{Normalized Weight}}\times\;\underbrace{\left(-\log\pi_{\theta}(a\mid o)\right)}_{\text{BC Loss}}.$ (2)
4 Experiments
Our aim in the experiments is to study the efficacy of HAND as a robot data retrieval pipeline and evaluate its ability to quickly learn to solve new downstream tasks. To this end, we organize our experiments to answer the following questions, in order:
@@ 4 Experiments
    How effective is HAND, using 2D relative paths, in retrieving task-relevant behaviors?
@@
    Does HAND work with hand demonstrations from unseen scenes?
@@
    Does HAND enable learning tasks in new scenes in simulation?
@@
    Can HAND enable real-time, fast adaptation on a real robot?
4.1 Experimental Setup
We evaluate HAND both in simulation using the CALVIN benchmark [8] and on real-world manipulation tasks with the WidowX-250 robot arm.
Figure 3: CALVIN Results. Task success rate of HAND and baseline methods on the CALVIN ABC-D task across three random seeds. Ablations of HAND are denoted by hatches. HAND and ablations outperform the next best baseline Flow on task success rate across all tasks.
CALVIN contains unstructured, teleoperated play data in four tabletop manipulation environments {A,B,C,D}, that share the same set of objects, but have different visual textures and static object locations (e.g., slider, button, switch), shown in Figure 6 (Left). Because it is infeasible to provide explicit human hand demonstrations in CALVIN, we instead perform end-effector point-tracking on expert task demonstrations to mimic the effect of hand-based tracking. We uniformly sample $N=6$ task-specific expert trajectories from environment D as $\mathcal{D}_{\text{hand}}$, and utilize about $17$k trajectories from environments {A,B,C} as $\mathcal{D}_{\text{play}}$. We evaluate our fine-tuned policy in environment D across 3 tasks.
Real World. We demonstrate that HAND can also scale to real-world scenarios by evaluating on several manipulation tasks in a kitchen setup shown in Figure 7. We collect a task-agnostic play dataset of about $50$k transitions. Human teleoperators were instructed to freely interact with the available objects in the scene without being bound to specific task goals. Object positions are randomized within the workspace during data collection and evaluation. We test three tasks: Reach Green Block, Press Button, and Close Microwave. We also introduce two additional difficult, long-horizon tasks, Put K-Cup in Coffee Machine and Blend Carrot, which require great precision and more than 150 real-world timesteps at a 5hz control frequency to execute, highlighting the capabilities of HAND to learn complex behaviors in real-time. Partial success is provided for tasks composed of multiple subtasks. Refer to Section A.2 for description of each task.
Baselines: We compare HAND to several retrieval baselines. All methods use the same transformer policy where applicable. We refer the reader to Figure 6 for implementation details and Appendix E for extensive ablation results. We consider the following baseline methods:
  • $\pi_{\text{base}}$ is the base policy pre-trained only on task-agnostic play data;
@@ 4.1 Experimental Setup
  • CLIP retrieves based on cosine similarity between target task language description’s CLIP embeddings (instead of hand demonstrations) and play data’s CLIP frame embeddings;
@@
  • Flow [11] trains a dataset-specific VAE on pre-computed optical flows for $\mathcal{D}_{\text{play}}$ from GMFlow [41] and retrieves individual states-action pairs based on latent motion similarity; and
@@
  • STRAP [12] also uses S-DTW for sub-trajectory retrieval but computes S-DTW distance based solely on Euclidean distance between pre-trained DINO-v2 image embeddings.
@@
STRAP and Flow assume access to expert robot demonstrations for both retrieval and fine-tuning. Unless otherwise stated, we adopt them for our setting without expert robot demonstration fine-tuning because we do not assume access to them. STRAP and Flow propose training policies from scratch, but we LoRA fine-tune both of them, like with HAND, because we found it to perform better.
4.2 Experimental Evaluation
Block Button Microwave
@@ 4.2 Experimental Evaluation
Table 1: Number of retrieved sub-trajectories performing demonstrated task. HAND retrieves more sub-trajectories performing the task compared to Flow and STRAP.
@@
(Q1): HAND retrieves more task-relevant data. We analyze the quality of retrieved sub-trajectories between Flow, STRAP  and HAND. STRAP and HAND both use S-DTW-based trajectory retrieval, but STRAP relies purely on visual DINO-v2 embeddings for retrieval. We provide a single hand demonstration of three real robot tasks and retrieve the top $K=25$ matches from $\mathcal{D}_{\text{play}}$. Compared to STRAP and Flow, we observe in Table 1 that HAND retrieves more trajectories in which the robot performs the demonstrated task. As both STRAP and Flow retrieve solely based on visual similarity, they perform poorly when there is a visual gap between the target demonstrations, e.g., human hand videos, and the offline robot play dataset. In particular, for the Push Button task, STRAP cannot retrieve any button pushing trajectories in its top 25 matches. Moreover, we ablate HAND’s visual filtering step and show that it helps retrieve $+30\%$ more relevant trajectories across all tasks. We provide qualitative comparisons of retrieved trajectories by in Appendix D.
@@
Figure 4: Qualitative retrieval results on out-of-distribution scene. We visualize the top sub-trajectory match of Flow , STRAP , HAND without visual filtering (HAND(-VF)), and HAND on two out-of-domain demonstrations recorded from an iPhone camera, showing approaching a K-Cup and putting it into the machine. Only HAND’s top match is relevant for both hand demonstrations.
@@
(Q2): HAND works with hand demos from unseen environments. Because HAND retrieves based on relative hand motions, it can work with target hand demos from out-of-distribution scenes, provided the camera angle remains relatively close to that in the play dataset. To demonstrate this scene robustness, we collect hand demos from a different scene with a handheld iPhone camera and a real coffee machine. We retrieve from robot play data containing a completely different scene and a toy coffee machine. In Figure 4, we show the lowest cost retrieved sub-trajectory of STRAP and Flow compared to HAND and a HAND ablation without the visual filtering step, HAND(-VF). Both of the retrieved trajectories for STRAP and Flow, along with the top trajectory for HAND(-VF) are irrelevant to the demonstrated task. Only HAND retrieves relevant robot trajectories for both hand demos because it focuses on the motion demonstrated by the human hand after visual filtering.
(Q3): HAND enables policy learning in simulation and real world. In Figure 3, we demonstrate that HAND and ablations outperform baselines in CALVIN, with a 16% average improvement over Flow and 123% over STRAP. In Figure 3, we also ablate the use of S-DTW-based loss weighting from Equation 2 with HAND(-CW), visual filtering from Equation 1 with HAND(-VF), and ground truth 3D pose information with HAND(+3D,-VF,-CW). HAND outperforms all of these ablations in Move Slider Left. Surprisingly, in this task, HAND(+3D,-VF,-CW) with priviledged 3D information, even underperforms HAND(-CW). We believe this is because, as HAND(+3D,-VF,-CW) retrieves trajectories based on an exact match in 3D end-effector pose, the retrieved trajectories have little variability and thus fail to generalize to changes in object placement in the scene. In some tasks, we notice that adding visual filtering can negatively impact performance, likely for a similar reason that filtering constrains the diversity of the resulting data subset. However, we demonstrated in the above two paragraphs that visual filtering helps in the real world to retrieve task-relevant trajectories.
@@
Figure 5: Real-Robot Results. Task completion out of 10 of $\pi_{\text{base}}$, STRAP, Flow, and HAND.
@@
Real-world experiments in Figure 5 demonstrate that fine-tuning with HAND improves success rates by +45% over the next best baseline, STRAP. In contrast, Flow fails to learn a policy that achieves reasonable success rates in any of the tasks, despite it being the best-performing baseline in CALVIN. Despite visual filtering not always helping in simulation in CALVIN, we observe that visual filtering is necessary in the real world to retrieve trajectories where the target object is interacted with, as demonstrated with HAND(-VF)’s worse retrieval performance in Table 1. We ablate different $K$ values for real robot tasks in Appendix F. We also report the performance of $\pi_{\text{base}}$, trained on all of $\mathcal{D}_{\text{play}}$.
@@
(Q4): HAND enables real-time, data-efficient policy learning of long-horizon tasks. We performed two small-scale user studies with IRB approval from our institution to demonstrate real-time learning. In the first study, a participant familiar with HAND iteratively demonstrated each part of a long-horizon Blend Carrot task (shown in Figure 7) and trained a HAND policy with over 70% success rate in under four (4) minutes from providing a single hand demonstration to deploying the fine-tuned policy. A video of a similar experiment can be found on our project website.
@@
Hand Demos (Min) $\downarrow$ 3 2
Robot Demos (Min) $\downarrow$ 10 14
Hand Demos (SR) $\uparrow$ 5/10 4/10
Robot Demos (SR) $\uparrow$ 3/10 2.5/10
@@
Table 2: Hand vs. Robot Teleoperation. Comparison of time taken and success rates between hand and teleoperated demonstrations.
@@
In the second study, two external users with prior teleoperation experience—but not affiliated with this research—each collected 10 demonstrations, using both their hand and robot teleoperation, to train the robot for the Put Keurig Cup in Coffee Machine task (see Figure 7). We employ HAND retrieval for hand-collected demonstrations and STRAP retrieval for robot teleoperation demonstrations. For a direct comparison, we additionally fine-tune STRAP with the human-collected teleoperated demonstrations as their paper suggests.
@@
As reported in Table 2, teleoperated demonstrations required over 3×3$\times$3 × more time to collect than hand demonstrations. Notably, using a single hand demonstration per user, we fine-tuned a policy exceeding 40% task completion compared to STRAP which only achieves 25% using a single robot teleoperation demonstration. We observed that increasing the number of expert demonstrations used for STRAP actually hurt downstream performance as it reduced the quality of retrieved trajectories. Thus, our results demonstrate that HAND is effective in enabling fast adaptation to downstream tasks.
5 Conclusion

### v4

@@ III-D Putting it All Together: Fast-Adaptation with Parameter-Efficient Policy Fine-tuning
Adapting to $\mathcal{D}_{\text{retrieved}}$. To rapidly adapt to a new task with minimal data, we leverage parameter-efficient fine-tuning using task-specific adapters—small trainable modules that modulate the behavior of the frozen base policy. Adapter-based methods have shown promise in few-shot imitation learning [40, 41], making them ideal for our limited retrieved dataset $\mathcal{D}_{\text{retrieved}}$. Specifically, we insert LoRA layers [42] into the transformer blocks of $\pi_{\text{base}}$. These are low-rank trainable matrices (about $0.1\%$$2\%$ of $\pi_{\text{base}}$’s parameters) inserted between the attention and feedforward layers (see Figure 2, LoRA Layers). During fine-tuning, we update only the parameters of these LoRA layers, $\theta$, using $\mathcal{D}_{\text{retrieved}}$.
Loss Re-Weighting. While our retrieval mechanism identifies sub-trajectories relevant to the target task, not all will be equally useful. Following prior work [14, 15, 25], we reweight the BC loss with an exponential term $\in(0,\infty)$ (similar to AWR [43]), where each sub-trajectory is weighted based on its S-DTW similarity to the hand demonstration. Intuitively, this upweights the loss of the most relevant examples in $\mathcal{D}_{\text{retrieved}}$ and downweights those that are less relevant. Finally, because trajectory cost scales vary depending on the task being retrieved and the features being used for S-DTW, we rescale the S-DTW costs $C_{i,\text{path}}$ to a fixed range. For each $\tau_{i}\in\mathcal{D}_{\text{retrieved}}$, its weight $e^{-C_{i,\text{path}}}$ is scaled to between $[0.01,100]$, where the normalization term comes from the sum of costs of all trajectories in $\mathcal{D}_{\text{retrieved}}$. Let the normalized weight for a trajectory be $w_{i}=\exp(-C_{i,\text{path}})$ and the behavioral cloning loss be $L_{i}(a,o)=-\log\pi_{\theta}(a\mid o)$. The total loss is then the weighted average over the dataset $\mathcal{D}$:
$\mathcal{L}_{\text{BC};\theta}=\frac{1}{|\mathcal{D}_{\text{retrieved}}|}\sum_{\tau_{i}\in\mathcal{D}_{\text{retrieved}}}w_{i}\times L_{i}(a,o).$ (2)
We summarize HAND in the pseudocode in Algorithm 1.
Algorithm 1 HAND Pseudocode
[…]
2:Train base policy $\pi_{\text{base}}$ on $\mathcal{D}_{\text{play}}$ via behavior cloning
3:Segment both $\mathcal{D}_{\text{play}}$ and $\mathcal{D}_{\text{hand}}$ into sub-trajectory datasets w/ threshold $\epsilon$: $\mathcal{T}_{\text{play}}$, $\mathcal{T}_{\text{hand}}$
4:for $\tau^{\text{hand}}\in\mathcal{T}_{\text{hand}}$ do
5:  Filter top-$M$ visually similar $\tau^{\text{play}}\in\mathcal{T}_{\text{play}}$ via DINO-based $\text{C}_{\text{visual}}$
6:  for each filtered $\tau^{\text{play}}$ do
7:   Track 2D hand paths with Molmo + CoTracker3
8:   Retrieve $K$ best-matching segments via S-DTW on relative path similarity
9:Fine-tune $\pi_{\text{base}}$ on retrieved data with adapter layers $\theta$ to obtain $\pi_{\theta}$ with $\mathcal{L}_{\text{BC};\theta}$ Equation 2
10:return $\pi_{\theta}$
IV Experiments
Our experiments demonstrate the efficacy of HAND as a robot data retrieval pipeline and evaluate its ability to quickly learn to solve downstream tasks. To this end, we organize our experiments to answer the following questions, in order:
@@ IV Experiments
    How well can HAND retrieve task-relevant behaviors?
@@
    Does HAND support hand demonstrations from unseen scenes and is it robust to visual shifts?
@@
    How does HAND perform in policy learning?
@@
    Can HAND enable real-time adaptation?
IV-A Experimental Setup
We evaluate HAND on a real-world multi-task kitchen environment using the WidowX robot arm. Our robot environment setup is shown in Figure 3. We use an Intel Realsense D435 camera as an external camera and a Logitech C920 as an over-the-shoulder camera.
Evaluation Tasks: We first evaluate on 10 total tasks. We first evaluate on three standard tasks: Reach Green Block, Press Button, and Close Microwave. Then, we introduce three challenging long-horizon tasks: Put K-Cup in Coffee Machine, Blend Carrot, and Cook Carrot, which demand high precision and span more than 150 timesteps at a 5 Hz control frequency. In particular, Cook Carrot is composed of four shorter tasks, Slide Pot $\rightarrow$ Put Object in Pot $\rightarrow$ Put Lid on Pot $\rightarrow$ Turn Stove Knob, including non-prehensile tasks (e.g., slide pot) and taking $\sim 300$ steps to complete even for expert teleoperators. For our long-horizon tasks, we provide one hand demonstration to perform retrieval for each subtask. These tasks highlight the ability of HAND to acquire and execute complex behaviors in real time. Partial success is provided for tasks composed of multiple subtasks.
Play Dataset Collection: We collect a task-agnostic play dataset containing a total of 50k transitions, each trajectory having an average of 230 timesteps and covering multiple tasks, collected at 5hz. The full dataset required roughly four hours to collect. We place distractor objects not used in target tasks in the environment that teleoperators interact with during play data collection to ensure the play data does not mirror evaluated tasks. The dataset is split into two, with about 1 hour corresponding to the scene for Cook Carrot. To evaluate language-conditioned methods, we manually annotate the Cook Carrot scene of the dataset with language, which takes an additional 87 minutes. During data collection and evaluation, movable task objects are randomized in a 5” x 7” region within the workspace.
Baselines: We compare HAND to the following baselines:
  • $\pi_{\text{base}}$: the base policy pre-trained on play data;
  • Lang Cond: $\pi_{\text{base}}$ with language-conditioning;
@@ IV-A Experimental Setup
  • CLIP (Language): retrieves based on cosine similarity between language embedding of target task (rather than hand demo) and language embedding of the play data;
  • CLIP (Image): retrieves based on cosine similarity between language embedding of target task and image embedding of the play data;
@@
  • Flow [12]: trains a VAE on pre-computed optical flows for $\mathcal{D}_{\text{play}}$ from GMFlow [44] and retrieves individual states-action pairs based on latent motion similarity; and
@@
  • STRAP [13]: also uses S-DTW for sub-trajectory retrieval but uses S-DTW distance based solely on Euclidean distance between pre-trained DINO-v2 image embeddings.
@@
STRAP and Flow assume access to expert robot demonstrations for both retrieval and fine-tuning. In our setting, we do not assume such demonstrations, therefore, unless otherwise noted, we adopt them without expert fine-tuning. While STRAP and Flow originally propose training policies from scratch, we instead apply LoRA fine-tuning—as with HAND —which we found to yield better performance for these baselines.
Policy Architecture: To ensure fair comparison, all methods use a three-layer action-chunking transformer (similar to ACT [36]) decoder policy where applicable. The input to the transformer policy is a sequence of image tokens corresponding to the external and over-shoulder camera views. Conditioned on the current image observation, the model predicts an action chunk corresponding to a second of execution.
Figure 3: WidowX Robot Arm Setup. We evaluate the scalability of HAND on 10 manipulation tasks on a WidowX robot arm in a kitchen setup [45].
Reach Green Block Push Button Close Microwave
@@
TABLE I: Number of retrieved sub-trajectories performing demonstrated task. HAND retrieves more task-performing sub-trajectories than Flow and STRAP.
IV-B Experimental Evaluation
@@
(Q1): HAND retrieves more task-relevant data. We analyze the quality of retrieved sub-trajectories between Flow, STRAP, and HAND. STRAP and HAND both use S-DTW-based trajectory retrieval, but STRAP relies purely on visual DINO-v2 embeddings for retrieval. We provide a single hand demonstration of three real robot tasks and retrieve the top $K=25$ matches from $\mathcal{D}_{\text{play}}$. Compared to STRAP and Flow, we observe in Table I that HAND retrieves more trajectories in which the robot performs the demonstrated task. STRAP relies exclusively on visual similarity, while Flow relies exclusively on motion similarity. Both methods struggle when there is a significant visual or motion gap between the target demonstrations (e.g., human hand videos) and the play dataset. In particular, for the Push Button task, STRAP is unable to retrieve any relevant trajectories in its top matches.
We also observe that visual filtering is necessary to retrieve trajectories where the target object is interacted with, as demonstrated by HAND(-VF), an ablation of HAND without visual filtering (Section III-C), having 30% worse retrieval performance than HAND in Table I.
@@
Figure 4: Qualitative retrieval results on OOD scene. We visualize the top sub-trajectory match of Flow, STRAP, HAND without visual filtering (HAND(-VF)), and HAND on two OOD domain demonstrations recorded from an iPhone camera, showing approaching a K-Cup and putting it into the machine. Only HAND’s top match is relevant for both hand demonstrations.
Method $10^{\circ}$ Horiz. $20^{\circ}$ Horiz. $30^{\circ}$ Horiz. $10^{\circ}$ Vert $20^{\circ}$ Vert $30^{\circ}$ Vert
Flow 0 / 25 2 / 25 5 / 25 1 / 25 0 / 25 0 / 25
STRAP 1 / 25 10 / 25 13 / 25 12 / 25 11 / 25 11 / 25
HAND 21 / 25 18 / 25 19 / 25 16 / 25 13 / 25 14 / 25
TABLE II: Camera angle robustness results. # of relevant retrieved trajectories for the Put Lid on Pot task if we change the camera angle vertically and horizontally by $10^{\circ}$ increments. HAND retrieves $18\%$ more relevant trajectories compared to STRAP even in the extreme case of $30^{\circ}$ shift.
@@
(Q2): HAND supports hand demonstrations from unseen environments and is robust to camera angle shifts. Because HAND retrieves based on relative hand motions, it is also effective with hand demonstrations from out-of-distribution (OOD) scenes. To illustrate, we collect hand demonstrations in a new environment using a handheld iPhone camera and a real coffee machine, while retrieving from robot play data recorded in a completely different scene with a toy coffee machine. In Figure 4, we show the lowest cost retrieved sub-trajectory of STRAP and Flow compared to HAND and a HAND ablation without the visual filtering step, HAND(-VF). Both of the retrieved trajectories for STRAP and Flow, along with the top trajectory for HAND(-VF) are irrelevant to the demonstrated task. For the first task, STRAP is able to retrieve the initial reaching motion toward the K-Cup but misses the crucial grasping segment, as it does not leverage motion for retrieval. Only HAND retrieves relevant robot trajectories for both hand demonstrations because it focuses on the motion demonstrated by the human hand after visual filtering.
Table II shows that HAND is also robust to shifts in camera angle for the Put Lid on Pot task, far more than Flow and STRAP. We measure the number of relevant retrieved trajectories of different methods after vertical and horizontal camera angle shifts of $10^{\circ}$ increments. In the most extreme setting of $30^{\circ}$ shift, HAND retrieves $18\%$ more relevant trajectories compared to STRAP. These camera angle shifts emulate head rotations on humanoid robots or camera movement on mobile manipulators, suggesting that HAND can even work in such settings where camera viewpoint change may occur.
@@
Figure 5: Real-Robot Results. Task completion (including partial completion) out of 10 of $\pi_{\text{base}}$, STRAP, Flow, and HAND.
Method
Slide
Pot
Put Obj
in Pot
Put Lid on
Pot
Turn
Knob
Long
Horiz.
Lang Cond 2 0 0 0 0
CLIP (Language) 0 4 1 0 0
STRAP 0 0 2 0 0
HAND w/o Pre-training 1 1 1 2 0
HAND(-VF) 2 0 0 5 0
HAND(-CW) 2 5 3 5 0
HAND 5 6 4 6 3
HAND + Lang Cond 8 7 5 7 5
TABLE III: Long horizon Cook Carrot task results. We show success rates on each subtask and on the full task execution. Full Task Success is out of 10.
Figure 6: Fast Adaptation Study. We conduct a small-scale user study to demonstrate HAND’s ability to learn robot policies in real-time. From providing the hand demonstration (Left), to retrieval and fine-tuning a base policy (Middle), to evaluating the policy (Right), we show HAND can learn to put a carrot in the blender with 7.5/10 task completion in less than 4 minutes.
@@
(Q3): HAND enables efficient policy learning in the real world. We evaluate four methods, including ours, across three standard tasks with ten trials each, for a total of 120 evaluations. Real-world experiments in Figure 5 on our three standard tasks demonstrate that fine-tuning with HAND improves success rates by +45% over the next best baseline, STRAP. In contrast, Flow fails to learn a policy that achieves reasonable success rates in any of the tasks. We also report the performance of $\pi_{\text{base}}$, trained on all of $\mathcal{D}_{\text{play}}$ and note that the pre-trained policy struggles to perform the tasks without any task-specific fine-tuning.
We next evaluate eight methods, including several ablations of HAND, across four base tasks and one long-horizon task constructed from these base tasks, with ten trials each for a total of 400 evaluations. Results on the more challenging long-horizon tasks (Table III) demonstrate that retrieval using hand demonstrations outperforms language-based retrieval (CLIP (Language)) by a factor of in success rate. Language-based retrieval suffers from the lack of spatial awareness, often retrieving trajectories that are semantically correct but spatially misaligned—similar to STRAP —which makes policy fine-tuning more difficult. In contrast, directly conditioning on language performs poorly compared to retrieval (Lang Cond), despite the fact that annotating sub-trajectories with language more than doubles data collection and annotation time.
Ablation Study: We observe that each component of HAND, namely pretraining, visual filtering (VF), and cost weighting (CW), are critical for task performance. Cost weighting helps bias the resulting policy towards behaviors that are most relevant to the downstream task, and reduces the effect of potentially noisy retrievals that may not directly aid in learning the target task. Without any of these components, the resulting policy is unable to learn the task. Only HAND successfully performs the entire long-horizon task reliably. We also show that given access to a language-annotated dataset, one could add language-conditioning on top of HAND to further improve the task performance (HAND + Lang Cond).
@@
(Q4): HAND enables real-time, data-efficient policy learning of long-horizon tasks. Finally, we performed two small-scale user studies with IRB approval from our institution to demonstrate real-time learning. In the first study shown in Figure 6, a participant familiar with HAND iteratively demonstrated each part of a long-horizon Blend Carrot task and trained a HAND policy with over 70% success rate in under four (4) minutes from providing a single hand demonstration to deploying the fine-tuned policy. A video of a similar experiment can be found on our website.
@@
Hand Demonstrations (Min) $\downarrow$ 3 2
Robot Demonstrations (Min) $\downarrow$ 10 14
Hand Demonstrations (SR) $\uparrow$ 5/10 4/10
Robot Demonstrations (SR) $\uparrow$ 3/10 2.5/10
@@
TABLE IV: Hand vs Robot Teleoperation. Time taken and success rates between hand and teleoperated demonstrations.
@@
Hand vs Robot Demonstration Comparison: In the second study, two users with prior teleoperation experience—but not affiliated with this research—each collected a total of 20 demonstrations: 10 using hand teleoperation and 10 using robot teleoperation, to train the robot for the Put K-Cup in Coffee Machine task. We employ HAND retrieval for hand-collected demonstrations and STRAP retrieval for robot teleoperation demonstrations. For a direct comparison, we additionally fine-tune STRAP with the human-collected teleoperated demonstrations as per [13].
@@
As reported in Table IV, teleoperated demonstrations required over $3\times$ more time to collect than hand demonstrations. Remarkably, with just a single hand demonstration per user, we fine-tuned a policy achieving over 40% task completion compared to STRAP which reaches only 25% using a single robot teleoperation demonstration. Interestingly, we observed that increasing the number of expert demonstrations used for STRAP degraded downstream performance likely due to lower quality retrieved trajectories. These results demonstrate that HAND enables fast adaptation to downstream tasks with as few as one (1) easy-to-provide hand demonstration.
V Conclusion and Limitations